OpenStudy (anonymous):

Can someone help me find the lateral area?? The slant height of a regular pentagonal pyramid is 8 cm and the length of each side of the base is 4 cm. What's the lateral area? a. 32 square cm b. 60 square cm c. 80 square cm d. 160 square cm I think it's C. Am I right?

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

so area AVB = 4*8/2 = 16 than multiplie this area by 5 and will get the lateral area what will be 5*16 = 80 cm^2 hope that you understand it sure

OpenStudy (anonymous):

thank you so much!